Solve the following problem


1. A gas occupies a volume of 500 cm 3 at 27⁰C. What will be the volume of the gas when

the temperature is -48⁰C? Assume pressure and moles are constant.


2. The pressure in an automobile tire is 29.8 psi on a cold morning of 20°C. After it is

driven a while, the temperature rises to 43°C. What is the pressure in the tire assuming

that the volume remains constant?





1
Expert's answer
2021-04-07T11:37:38-0400

Solution:

Question 1:

T1 = 27°C + 273 = 300 K

V1 = 500 cm3

T2 = -48°C + 273 = 225 K

V2 = unknown

P1 = P2 = const

n1 = n2 = const


Since the pressure and amount of gas remain unchanged, Charles' Law can be used.

Charles' Law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the volume, solve the equation for V2:

V2 = V1T/ T1

V2 = (500 cm3 × 225 K) / (300 K) = 375 cm3

V2 = 375 cm3

 

Answer to Question 1: The volume of the gas will be 375 cm3.


Question 2:

P1 = 29.8 psi 

T1 = 20°C + 273 = 293 K

P2 = unknown

T2 = 43°C + 273 = 316 K

V1 = V2 = const

n1 = n2 = const


Since the volume and amount of gas remain unchanged, Gay-Lussac's Law can be used.

Gay-Lussac's Law can be expressed as: P1 / T1 = P2 / T2

P1T2 = P2T1

To find the pressure, solve the equation for P2:

P2 = P1T2 / T1

P2 = (29.8 psi × 316 K) / (293 K) = 32.14 psi

P2 = 32.14 psi


Answer to Question 2: The pressure in the tire will be 32.14 psi.